The average annual number of cigarettes smoked by an adult in some country continues to decline. For the years 2000-2012 the equation y=-77.8x+2051 approximates this data.​ Here, x is the number of years after​ 2000, and y is the average annual number of cigarettes smoked. If this trend​ continues, find the year in which the average annual number of cigarettes smoked reaches zero. To do​ this, substitute 0 for y in the equation and solve for x.


1
Expert's answer
2022-01-19T05:55:08-0500

Y = -77.8 X + 2051

X = No. of years

Y = average annual number of cigarettes smoked

Put Y = 0 (Target Value of Y)

Therefore,

0 = -77.8 X + 2051

Or     77.8 X = 2051

Or     X = 2051/77.8

Or     X = 26.4


Final Year = 2000 + X = 2000 + 26 = 2026